Ana Carillo and Associates is a medium-sized company located near a large metropolitan area in the Midwest. The company manufactures cabinets of mahogany, oak, and other fine woods for use in expensive homes, restaurants, and hotels. Although some of the work is custom, many of the cabinets are a standard size.

One such non-custom model is called Luxury Base Frame. Normal production is 1,000 units. Each unit has a direct labor hour standard of 5 hours. Overhead is applied to production based on standard direct labor hours. During the most recent month, only 950 units were produced; 4,500 direct labor hours were allowed for standard production, but only 4,000 hours were used. Standard and actual overhead costs were as follows.

Standard
(1,000 units)
Actual
(950 units)
Indirect materials $  12,800 $  13,100
Indirect labor 45,700 54,300
(Fixed) Manufacturing supervisors salaries 23,900 23,400
(Fixed) Manufacturing office employees salaries 13,800 13,300
(Fixed) Engineering costs 28,700 26,600
Computer costs 10,600 10,600
Electricity 2,700 2,700
(Fixed) Manufacturing building depreciation 8,500 8,500
(Fixed) Machinery depreciation 3,200 3,200
(Fixed) Trucks and forklift depreciation 1,600 1,600
Small tools 700 1,500
(Fixed) Insurance 500 500
(Fixed) Property taxes 300 300
  Total $153,000 $159,600

Calculate the total overhead variance, controllable variance, and volume variance. (Round variable overhead to 2 decimal places and final answers to 0 decimal places, e.g. 1,575.)

Total overhead variance $

Entry field with correct answer

Entry field with correct answer UnfavorableNeither favorable nor unfavorableFavorable

Controllable variance $

Entry field with incorrect answer now contains modified data

Entry field with correct answer UnfavorableNeither favorable nor unfavorableFavorable

Volume variance $

Entry field with incorrect answer now contains modified data

Entry field with correct answer FavorableUnfavorableNeither favorable nor unfavorable
Answer #1

* overhead opplication rate = Total Standard overhead / Total Standard hours 153000 / (looox 5) 30.6* Appleed overhead - standard hours for actual production X capplecation rate = Overhead 4500 x 30.6 137700Total overhead varlance = = = Appleed overhead - Actual overhead 137700 - 159600 21900 WTotal budgeted variable overhead = 19800 +45700 + 10600 +2700 + 700 = 72500 Variable overhead rate per hour = 12500 / 5000 =